I feel like a dumb parent as I cant help my child with this. please help.
Cell Company offers a single line plan for $50. they wan to offer a family plan for 2 lines or more under the same plan. doing so you will decrease the charges per line by $5 each, (the costs per line would decrease for all lines, not just for the lines beyond the first 2.
1. What is the costs per line?
2. If you have five lines under the same family plan what will the total costs be.
3. if you have 8 lines under the same family plan what will be the total costs.
4. what would be the factored from and the general form for the total costs of the family plan as a function of the number of lines beyond 2

Answer by ankor@dixie-net.com(22740)

Cell Company offers a single line plan for $50.
they want to offer a family plan for 2 lines or more under the same plan. doing so you will decrease the charges per line by $5 each, (the costs per line would decrease for all lines, not just for the lines beyond the first 2.
:
I am assuming that each line is $5 less than the previous line
:
Let x = no. of lines
1. What is the costs per line?
50 - 5(x-1)
:
2. If you have five lines under the same family plan what will the total costs be.
x(50 - 5(x-1))
x=5
5(50 - 5(5-1)) = 5(50-5(4)) = 5(50-20) = 5(30) = $150 for five lines
:
3. if you have 8 lines under the same family plan what will be the total costs.
x=8
8(50 - 5(8-1)) = 8(50 - 35) = 8(15) = $120 for eight lines
:
4. what would be the factored form and the general form for the total costs of the family plan as a function of the number of lines beyond 2
f(x) = x(50 - 5(x-1))
f(x) = x(50 - 5x + 5)
f(x = 50x - 5x^2 + 5x
f(x) = -5x^2 + 55x
can be factored to
f(x) = -5x(x - 11)
This is interesting in that if you get 11 lines, you pay 0!
If you get 12 lines they pay you $60!
Maybe I missed something here
:
:
You are not a dumb parent, you are a smart parent, the dumb parent does not try to personally help his child. CK